OpenStudy (anonymous):

A microwave oven operates at a frequency of approximately 2450 MHz. What is the corresponding wavelength?

OpenStudy (anonymous):

Well then, take note that all electromagnetic wave have the speed of light, which is 3*10^8 m/s. Then use the speed equation for waves: v=f*lambda, where v is speed, f is frequency and lambda is wavelength. Replace your values and fing for lambda. Do not forget to convert your frequency value to Hz. Do not leave it as MHz. That is the same as multiplying your 2450 by one million at first.

OpenStudy (mayankdevnani):

\[\large \bf frequncy(\nu)=2450~MHz=2450 \times 10^6Hz\]

OpenStudy (mayankdevnani):

as we know the relation :- \[\large \bf Velocity(v)=frequency(\nu) \times wavelength(\lambda)\] where,\[\large \bf velocity=3 \times 10^8m/s=speed~of~light(as~a~reference)\]

OpenStudy (mayankdevnani):

plug the values, \[\large \bf 3 \times 10^8=2450 \times 10^6 \times \lambda\] solve for lambda ?

OpenStudy (mayankdevnani):

@mrrodriguez can you solve it ?

OpenStudy (anonymous):

I can definitely attempt :)

OpenStudy (mayankdevnani):

good !! now tell me what you get lambda ?

OpenStudy (anonymous):

-2.35x10^7? @mayankdevnani

OpenStudy (mayankdevnani):

no,how did you get `-`sign ?

OpenStudy (mayankdevnani):

here is the solution :- \[\large \bf \lambda=\frac{3 \times 10^8}{2450 \times 10^6}=0.0012 \times 10^2=0.12m\]

OpenStudy (mayankdevnani):

hope you understand. @mrrodriguez

OpenStudy (anonymous):

That makes a lot more sense...I just didn't set the problem up right

OpenStudy (mayankdevnani):

don't worry !!

OpenStudy (mayankdevnani):

but i am happy that you understand !

OpenStudy (anonymous):

Thank you for your help....I hope you'll be on for a while cause I have many more questions :)

OpenStudy (mayankdevnani):

sorry :(..i have to go now !! for lunch ! my mouth is watering...not wasting so much time. i have to go..bbbyee. and sorry :(

OpenStudy (mayankdevnani):

i hope you will learn from @Hoslos

OpenStudy (anonymous):

I think I could help @mrrodriguez.

OpenStudy (mayankdevnani):

he is a great tutor of CHEMISTRY !

OpenStudy (anonymous):

Any help is much appreciated cause Chemistry is not my forte!

OpenStudy (mayankdevnani):

correct !! my friend

OpenStudy (anonymous):

Hehe, Great tutor? I doubt it, but thanks for the complement.